# Regularized Conditional Diffusion Model for Multi-Task Preference Alignment

### v1

@@ F.3 What enables the generalization to unseen tasks?
Appendix G Acceleration Sampling for Diffusion models
While multi-step denoising process in the diffusion model generation is time-consuming, our contribution is orthogonal to those sampling acceleration methods for diffusion models, such as DDIM Song et al. [2020], DPM-solver Lu et al. [2022], EDP Kang et al. [2024], and can be easily combined with them. In fact, we have incorporated the implementation of the DPM-solver into our method.. This improvement has boosted the inference speed by a factor of 8.4 compared to the previous method. Specifically, on an Nvidia RTX3090, the denoising time per step has been reduced from 15.2ms to 1.8ms. However, our initial results indicate a compromise in the quality of the generated trajectories, likely due to the complexity of generating continuous trajectories compared to image-generation tasks. We will explore further enhancements to balance performance and quality in our future work.
Appendix H Further Analysis About MI Regularization
The conditional diffusion model is not enough to provide alignment
The remarkable success of Stable Diffusion and Midjourney has highlighted the potential of diffusion models to generate desired images given textual conditions. However, when diffusion models are applied to offline reinforcement learning and trajectory modeling, the situation appears different. Previous work on return-conditioned diffusion models Ajay et al. [2023], Yuan et al. [2024] generates trajectories by conditioning on target returns, but their conditions are often hyperparameters adjusted for each environment. As shown in Figure 1, our tests with different conditions reveal that the relationship between the generation and the conditions is not as expected. Given a higher target return, the diffusion model cannot generate better trajectories. In fact, this phenomenon has been noted in other related works as well Tai et al. [2023]. Therefore, the current conditional diffusion models are insufficient to provide the necessary alignment between preference and trajectory in our setting. This is one of the main problems we aim to solve.
Why using MI regularization?
Inspired by the work in the field of image generation, such as InfoVAE Zhao et al. [2017], InfoGAN Chen et al. [2016], and InfoDiffusion Wang et al. [2023a], we propose the adoption of mutual information regularization. During the learning phase of the diffusion model, it effectively estimates prior noise, with its posterior estimation conditioned on preference representation. In our debugging phase, we observed that noise estimates from previous works often disregard conditioning, leading to the diffusion model’s inability to effectively learn the conditional distribution $p(\tau|c)$. Therefore, in our work, we imposed a constraint of maximizing mutual information on the diffusion model, ensuring a tighter relationship between the posterior noise estimation and the condition. This guarantees that the noise estimation network can capture information about the condition. Consequently, during the denoising generation phase, the trajectories generated by the diffusion model can well correspond to the condition. Moreover, in response to the inefficiency of the multi-step denoising process in diffusion models, we have introduced a reasonable approximation that simplifies the implementation of regularization. The experimental results also demonstrate that enhancing this connection leads to better outcomes than not doing so.

### v2

@@ H.3 What enables the generalization to unseen tasks?
Appendix I Space and Time Complexity
The computation in our method primarily involves two parts: 1) Training the MI Regularized Diffusion Model: Estimating the mutual information between the generated trajectory $\tau_{0}$ and the preference representation $w$ is computationally intensive, as generating $\tau_{0}$ requires $k$-step denoising. To mitigate this, we apply approximations and replace $I(\tau_{0};w)$ with $I(\tau_{k};w)$, thus bypassing the denosing process during training. $I(\tau_{k};w)$ can be obtained from the estimated noise at each step, requiring only a two-layer MLP to align the dimensions with $w$. 2) Inference Stage: Sampling during the inference stage is time-consuming, which we have discussed in Appendix F. To address this, we have experimented with DPM-solver [1] to reduce the sampling time of the diffusion model. This improvement has boosted the inference speed by a factor of 8.4 compared to the previous method. Specifically, on an Nvidia RTX3090, the denoising time per step has been reduced from 15.2ms to 1.8ms. Despite the improvement in efficiency, our initial results indicate a compromise in the quality of the generated trajectories, likely due to the complexity of generating continuous trajectories. We will explore further enhancements to balance efficiency and quality in our future work.
We have also conducted a further comparison of the space complexity and time complexity of CAMP, with the diffusion model-based multitask learning method, MTDiff. According to the results in Table 7, the runtime for each update in MTDiff is approximately half of that in CAMP. This is because, in addition to optimizing the diffusion model, CAMP also optimizes the preference representation network and the optimal preference representation. It is worth noting that the learning of preference representation and the optimization of the diffusion model can be decoupled; if decoupled, our method might exhibit greater flexibility and efficiency. On the other hand, by comparing the GPU memory usage during algorithm execution, we found that the space required by CAMP is about one-third of that required by MTDiff. This indicates that, although CAMP is more computationally intensive, its memory space requirements are significantly lower. This reason may be that CAMP utilizes a U-Net based noise predictor, while MTDiff needs a more complex Transformer-based noise predictor.
Table 7: Comparison of MTDiff and CAMP
MTDiff CAMP
Runtime per update (s) 32.0 62.5
Memory used (MB) 7322 2516
Appendix J Broader Impacts
This paper presents a novel approach to sequential decision-making, aiming to advance the field of machine learning. Our method addresses critical challenges in aligning decision-making processes with human intents and achieving versatility across various tasks. By adopting multi-task preferences as a unified condition for decision-making, we offer a framework that enhances both alignments with human preferences and versatility across tasks.
The societal consequences of our work are manifold. By improving the controllability of decision-making processes concerning human preferences, our approach holds promise for applications in diverse domains such as robotics, healthcare, and personalized recommendation systems. Additionally, by addressing limitations in existing methods regarding the formulation of reward functions, we pave the way for more robust and adaptable machine learning systems.
Further, we believe our work is beneficial for the broader research community in preference learning. For instance, in the alignment of large language models, researchers are increasingly recognizing the importance of diverse human preferences for multi-objective preferences[61, 62, 63, 64, 65]. Our work actually offers a perspective on this issue. For text-to-image or text-to-video generation, better alignment with human preferences is also a critical consideration[66], and our work holds relevant implications for these fields.